# Generalized Correlation Regression for Disentangling Dependence in Clustered Data

### v1

@@ Abstract
Clustered and longitudinal data are pervasive in scientific studies, from prenatal health programs to clinical trials and public health surveillance. Such data often involve non-Gaussian responses—including binary, categorical, and count outcomes—that exhibit complex correlation structures driven by multilevel clustering, covariates, over-dispersion, or zero inflation. Conventional approaches such as mixed-effects models and generalized estimating equations (GEEs) can capture some of these dependencies, but they are often too rigid or impose restrictive assumptions that limit interpretability and predictive performance.
@@
Keywords: Clustered data analysis; Correlation regression; Generalized $z$-transformation; Longitudinal data; Regression modeling
@@ 1.1 Background and Overview
As detailed in Section 3, our motivating example is the modern prenatal care dataset, which records clustered binary outcomes from a study conducted by researchers at Princeton University (Pebley et al.,, 1996; Rodríguez and Goldman,, 2001) to examine the effects of community-, maternal-, and individual-level covariates on prenatal care utilization. The data originate from the 1987 National Survey of Maternal and Child Health in Guatemala and, after preprocessing, consist of 2,223 births from 1,414 mothers across 160 communities. Covariates are available at three levels: community-level (e.g., accessibility of healthcare facilities), maternal-level (e.g., education and socioeconomic status), and individual-level (e.g., maternal age), creating a multi-level clustering structure with correlations induced both within mothers and within communities.
The binary response indicates whether a given birth received modern prenatal care. Conventional models can explain such choices through community-, maternal-, and individual-level covariates, but they struggle with the complex dependence structure inherent in the data. Responses are not independent: births from the same mother and mothers from the same community share unobserved factors that induce correlation, as recognized in prior studies. Crucially, our analysis reveals that dependence extends beyond hierarchical clustering. Correlations in prenatal care choices also emerge across mothers with similar covariates, for instance comparable socioeconomic or geographic contexts, leading to shared patterns of healthcare access. From a substantive perspective, detecting such dependence is critical: it highlights hidden sources of inequity and points to areas where targeted field studies may deliver the most effective interventions.
This example illustrates three key statistical challenges:
@@
    Complex within-cluster dependence, induced by multiple hierarchical levels (births nested within mothers, mothers nested within communities);
@@
    Non-Gaussian outcomes, where binary responses preclude direct application of Gaussian-based correlation modeling;
@@
    The need for flexible, interpretable correlation structures that capture both traditional clustering effects and dependence driven by covariates.
Motivated by these challenges, the main scientific question we address can be informally summarized as a regression model for correlations:
$\text{(Informal)}\quad\text{Correlation matrix }{\bf R}\;\sim\;\text{Covariates},$
where the covariates may include both continuous and categorical variables, drawn from multiple hierarchical levels. Unlike conventional regression, which links covariates to the mean response, our framework links them directly to the dependence structure, enabling a flexible yet interpretable characterization of within- and cross-cluster correlations.
@@ 1.3 Our Contributions
The paper is organized as follows. Section 2 presents the proposed methodology, including model specification, the generalized $z$-transformation, and parameter estimation with the computational algorithm. Section 3 illustrates the approach using the prenatal care dataset. Additional technical details, simulations, and two further applications—a toenail infection trial and clustered health count data—are in the Supplementary Material. All code is publicly available on https://github.com/yibowang2004/gcr for reproducibility and practical use.
@@ 2.2 Our Approach with Generalized z-Transformation
We begin by introducing some notation. Let $\mathbf{A}\in\mathbb{R}^{m\times m}$ be a symmetric matrix. The operator $\operatorname{vecl}(\mathbf{A})$ denotes the vector obtained by stacking the strictly lower-triangular elements of $\mathbf{A}$ in column-major order, so that $\operatorname{vecl}(\mathbf{A})\in\mathbb{R}^{m(m-1)/2}$. The operator $\operatorname{diag}$ has two distinct roles: (i) When applied to a vector $\mathbf{v}=(v_{1},\ldots,v_{m})^{\top}$, it produces a diagonal matrix: $\operatorname{diag}(\mathbf{v})=\mathrm{diag}(v_{1},\ldots,v_{m})$. (ii) When applied to a square matrix $\mathbf{A}$, it extracts the diagonal elements as a vector: $\operatorname{diag}(\mathbf{A})=(a_{11},a_{22},\ldots,a_{mm})^{\top}\in\mathbb{R}^{m}$. For a symmetric positive definite matrix $\mathbf{A}$ with spectral decomposition $\mathbf{A}=\mathbf{Q}\,\mathrm{diag}(\lambda_{1},\ldots,\lambda_{m})\,\mathbf{Q}^{\top}$, where $\mathbf{Q}$ is orthogonal and $\lambda_{j}>0$, the matrix exponential and logarithm are defined as
$e^{\mathbf{A}}=\mathbf{Q}\,\mathrm{diag}(e^{\lambda_{1}},\ldots,e^{\lambda_{m}})\,\mathbf{Q}^{\top},\qquad\log\mathbf{A}=\mathbf{Q}\,\mathrm{diag}(\log\lambda_{1},\ldots,\log\lambda_{m})\,\mathbf{Q}^{\top}.$
@@
where $\boldsymbol{\alpha}\in\mathbb{R}^{d}$ is the matrix log-correlation parameter and ${\mathbf{w}}_{ijk}$ contains pair-specific covariates capturing factors that explain the correlation between observations $j$ and $k$. The covariates ${\mathbf{w}}_{ijk}$ are flexible and may include:
  • differences or distances between ${\mathbf{x}}_{ij}$ and ${\mathbf{x}}_{ik}$,
  • subgroup indicators (e.g., same higher-level cluster),
  • other contextual or structural features relevant to the study.
Concrete constructions are illustrated in Section 3.
@@
initialized from an arbitrary $\mathbf{x}_{(0)}$.
This provides a constructive algorithm to recover a valid correlation matrix ${\bf R}=f^{-1}(\boldsymbol{\gamma})$ from an unconstrained vector $\boldsymbol{\gamma}$:
  1. 1.
    Embed $\boldsymbol{\gamma}$ into a symmetric matrix ${\bf G}=\operatorname{vecl}^{-1}(\boldsymbol{\gamma})$ with arbitrary diagonal entries;
  2. 2.
    Compute the unique diagonal adjustment $\mathbf{x}^{*}$ using the fixed-point iteration;
  3. 3.
    Obtain the correlation matrix as ${\bf R}=e^{{\bf G}[\mathbf{x}^{*}]}$.
2.3 Estimation and Algorithm
@@
where $C$ is a constant independent of $\boldsymbol{\alpha}$ and $\hat{{\bf V}}_{i}=\hat{{\bf A}}_{i}^{1/2}{\bf R}_{i}(\boldsymbol{\alpha})\hat{{\bf A}}_{i}^{1/2}$.
Defining the standardized residual $\hat{\boldsymbol{\nu}}_{i}=\hat{{\bf A}}_{i}^{-1/2}({\mathbf{y}}_{i}-\hat{\boldsymbol{\mu}}_{i}),$ the objective function simplifies to
$\sum_{i=1}^{n}\Big{\{}\log|{\bf R}_{i}(\boldsymbol{\alpha})|+\hat{\boldsymbol{\nu}}_{i}^{\top}{\bf R}_{i}(\boldsymbol{\alpha})^{-1}\hat{\boldsymbol{\nu}}_{i}\Big{\}}.$
Consequently, the pseudo-likelihood estimator of $\boldsymbol{\alpha}_{0}$ can be written as
@@
This approach offers a practical, numerically stable way to estimate correlation parameters in the generalized correlation regression framework. By working in the unconstrained space of the generalized $z$-transformation, it ensures valid correlation matrices while flexibly incorporating covariates. Under standard regularity conditions, the estimator $\hat{\boldsymbol{\alpha}}_{n}$ is consistent for $\boldsymbol{\alpha}_{0}$ and asymptotically normal, providing a sound basis for statistical inference; see the Supplementary Material for technical details.
To jointly estimate $(\boldsymbol{\alpha},\boldsymbol{\beta},\phi)$, we employ an iterative procedure based on a modified Fisher scoring algorithm that alternates between updates of $\boldsymbol{\beta}$, $\boldsymbol{\alpha}$, and $\phi$, combining the score functions from the GEE (4) with the pseudo-likelihood (9). Define the score functions as
${\bf A}=({\bf Q}\otimes{\bf Q})\boldsymbol{\Xi}({\bf Q}\otimes{\bf Q})^{\top},$
where $\boldsymbol{\Xi}$ is diagonal with entries
Algorithm 1 Modified Fisher’s Scoring Algorithm
0: Estimates $(\boldsymbol{\alpha},\boldsymbol{\beta},\phi)$
1:repeat
2:  Update $\phi^{(k)}$ using $\boldsymbol{\beta}^{(k-1)}$ as per (5) and (6)
4:  repeat
5:   Update
6:   Increment $s\leftarrow s+1$
7:  until convergence; update $\boldsymbol{\alpha}^{(k)}$
8:  Update
9:  Increment $k\leftarrow k+1$
10:until convergence
High-dimensional $\boldsymbol{\eta}_{i}$ can slow convergence. To mitigate this, we replace ${\bf J}_{i}$ with a pseudo-expectation $\tilde{\operatorname{\mathbb{E}}}({\bf J}_{i})$ capturing the dominant first-order structure while ignoring higher-order dependencies (see Supplementary Material), yielding stable estimates with negligible bias at lower cost.
@@ 3.1 Dataset Overview
3.2 Baseline Clustering Effects
A primary source of dependence arises from shared community- and family-level influences, expected to induce correlations among prenatal care choices. To visualize these baseline clustering effects, we examine the residuals
$\tilde{\varepsilon}_{ij}=\frac{y_{ij}-\hat{p}_{ij}}{\sqrt{\hat{p}_{ij}(1-\hat{p}_{ij})}},$
from fitting the mean model (3.1) via GEE with an independent working correlation. Empirical correlations within predefined subgroups $S$ are then computed as
$\hat{\rho}_{S}=\frac{1}{N_{S}}\sum_{(i,j)\neq(t,s)}\tilde{\varepsilon}_{ij}\,\tilde{\varepsilon}_{ts}\,\mathbb{I}\big{(}(y_{ij},y_{ts})\in S\big{)},$
where $i,t$ index communities, $j,s$ index records, and $N_{S}$ is the number of observation pairs in subgroup $S$. At the community level, let $S_{1}$ denote pairs within the same community ($i=t$ and $(i,j)\neq(t,s)$) and $S_{2}$ its complement. At the family level, $S_{3}$ (within-family) and $S_{4}$ (between-family) are defined analogously.
The resulting empirical correlations are
$\hat{\rho}_{S_{1}}=0.1455,\quad\hat{\rho}_{S_{2}}=-0.0015,\qquad\hat{\rho}_{S_{3}}=0.7764,\quad\hat{\rho}_{S_{4}}=-0.0007.$
Figure 1 presents boxplots of the residuals for the different subgroups, illustrating the presence of strong within-family correlation and moderate within-community correlation.
@@ 3.2 Baseline Clustering Effects
These results indicate that correlations are stronger within communities and within families, with family-level clustering effects notably more pronounced than community-level effects.
@@
where
@@
In this parametrization, $\alpha_{0}$ captures the community-level clustering effect, while $\alpha_{1}$ measures the family-level clustering effect.
@@
The estimated parameters are $\hat{\alpha}_{0}=0.0468$ and $\hat{\alpha}_{1}=0.8617$, both highly significant (p-values <0.0001<0.0001). These results are consistent with the empirical residual correlations and prior findings (Pebley et al.,, 1996; Rodríguez and Goldman,, 2001), confirming strong family-level and moderate community-level clustering effects in the data.

### v2

@@ Abstract
Clustered and longitudinal data are pervasive in scientific studies, from prenatal health programs to clinical trials and public health surveillance. Such data often involve non-Gaussian responses—including binary, categorical, and count outcomes—that exhibit complex correlation structures driven by multilevel clustering, covariates, over-dispersion, or zero inflation. Conventional approaches such as generalized mixed-effects models (GLMMs) and generalized estimating equations (GEEs) can capture some of these dependencies, but they are often too rigid or impose restrictive assumptions that limit interpretability and predictive performance.
@@
Keywords: Clustered data; Longitudinal data; Correlation regression; Generalized $z$-transformation; Regression modeling
@@ 1.1 Background and Overview
As detailed in Section 3, our primary illustrating example is the modern prenatal care dataset, a landmark study conducted by researchers at Princeton University (Pebley et al.,, 1996; Rodríguez and Goldman,, 2001). The dataset records clustered binary outcomes on whether births received modern prenatal care, drawn from the 1987 National Survey of Maternal and Child Health in Guatemala. After preprocessing, it consists of 2,223 births from 1,414 mothers across 160 communities, with covariates measured at three levels: community-level (e.g., accessibility of healthcare facilities), maternal-level (e.g., ethnic-linguistic identity and socioeconomic status), and individual-level (e.g., maternal age). This creates a natural multi-level clustering structure, where correlations arise both within mothers and within communities.
This dataset raises several critical scientific questions of both methodological and applied importance:
@@
    Covariate effects: How do community-, maternal- and individual-level covariates shape prenatal care utilization?
@@
    Dependence beyond hierarchy: Are there systematic correlations in care choices across mothers with similar socioeconomic or geographic backgrounds?
@@
    Equity and intervention: How can identifying such dependence patterns highlight sources of inequity and inform targeted, cost-effective public health interventions?
Answering these questions is of profound practical importance. Prenatal care utilization is widely recognized as a key determinant of maternal and infant health outcomes, and identifying the drivers of correlation—whether arising from community, maternal, or cross-level factors—has direct implications for targeting interventions. Detecting such dependence not only clarifies sources of inequity but also informs the design of cost-effective policies and field studies aimed at improving maternal and child health in resource-constrained settings.
The binary response indicates whether a given birth received modern prenatal care. Existing analyses typically model such choices using community-, maternal-, and individual-level covariates, but they often ignore or only partially capture the dependence inherent in the data. In reality, responses are not independent: births from the same mother and mothers from the same community share unobserved influences that induce correlation. Our analysis further reveals that dependence extends beyond hierarchical clustering, with correlations also arising across mothers who share similar covariates—for example, those in comparable socioeconomic contexts. Such patterns indicate potential systematic inequities in healthcare access that conventional models overlook. From a substantive perspective, detecting these cross-cluster dependencies is essential: it confirms the role of observed covariates, reveals hidden sources of inequity, and identifies population subgroups where targeted field studies and interventions may be most effective.
This example highlights three central statistical challenges:
  1. 1.
    Complex within-cluster dependence, arising from multiple hierarchical levels (births nested within mothers, and mothers nested within communities);
  2. 2.
    Non-Gaussian outcomes, where binary responses preclude direct use of Gaussian-based correlation modeling;
  3. 3.
    The need for flexible and interpretable correlation structures that capture not only traditional clustering effects but also dependence driven by covariates.
Motivated by these challenges, we propose an approach that can be informally summarized as a regression model for correlations:
$\text{(Informal)}\qquad\text{Correlation matrix }{\bf R}\;\sim\;\text{Covariates},$
where the covariates may include both continuous and categorical factors, drawn from multiple hierarchical levels. Unlike conventional regression, which links covariates to the mean response, our framework links them directly to the dependence structure, enabling a flexible yet interpretable characterization of within- and cross-cluster correlations.
The modeling device is key to answering the scientific questions raised by our motivating datasets. By treating correlations as regression targets, it uncovers dependence patterns beyond mean-based analyses. In the prenatal care study, for instance, it enables formal testing of whether correlations in healthcare choices arise not only within mothers or communities but also across mothers with similar socioeconomic profiles, thereby uncovering potential inequities that conventional models fail to detect, highlighting questions worth further field study or experimental investigation.
@@ 1.3 Our Contributions
The paper is organized as follows. Section 2 presents the proposed methodology, including model specification, the generalized $z$-transformation, and parameter estimation. Section 3 illustrates the approach using the prenatal care dataset. Section 4 provides an additional example on a longitudinal binary dataset. Additional technical details, simulations, and further data analysis are in the Supplementary Material. All code is publicly available on https://github.com/yibowang2004/gcr for reproducibility and practical use. All GEE and GLMM analyses in this paper are conducted using the gee package (Carey,, 2024) and the glmmTMB package (McGillycuddy et al.,, 2025) in R, respectively.
@@ 2.2 Our Approach with Generalized z-Transformation
We begin by introducing some notation. Let $\mathbf{A}\in\mathbb{R}^{m\times m}$ be a symmetric matrix. The operator $\operatorname{vecl}(\mathbf{A})$ denotes the vector obtained by stacking the strictly lower-triangular elements of $\mathbf{A}$ in column-major order, so that $\operatorname{vecl}(\mathbf{A})\in\mathbb{R}^{m(m-1)/2}$. The operator $\operatorname{diag}$ has two distinct roles: (i) When applied to a vector $\mathbf{v}=(v_{1},\ldots,v_{m})^{\top}$, it produces a diagonal matrix: $\operatorname{diag}(\mathbf{v})=\mathrm{diag}(v_{1},\ldots,v_{m})$. (ii) When applied to a square matrix $\mathbf{A}$, it extracts the diagonal elements as a vector: $\operatorname{diag}(\mathbf{A})=(a_{11},a_{22},\ldots,a_{mm})^{\top}\in\mathbb{R}^{m}$. For a symmetric positive definite matrix $\mathbf{A}$ with spectral decomposition $\mathbf{A}=\mathbf{Q}\,\mathrm{diag}(\lambda_{1},\ldots,\lambda_{m})\,\mathbf{Q}^{\top}$, where $\mathbf{Q}$ is orthogonal and $\lambda_{j}>0$, the matrix exponential and logarithm are defined as $e^{\mathbf{A}}=\mathbf{Q}\,\mathrm{diag}(e^{\lambda_{1}},\ldots,e^{\lambda_{m}})\,\mathbf{Q}^{\top}$ and $\log\mathbf{A}=\mathbf{Q}\,\mathrm{diag}(\log\lambda_{1},\ldots,\log\lambda_{m})\,\mathbf{Q}^{\top}$.
@@
where $\boldsymbol{\alpha}\in\mathbb{R}^{d}$ is the matrix log-correlation parameter and ${\mathbf{w}}_{ijk}$ contains pair-specific covariates capturing factors that explain the correlation between observations $j$ and $k$. The covariates ${\mathbf{w}}_{ijk}$ are flexible and may include: (i) differences or distances between ${\mathbf{x}}_{ij}$ and ${\mathbf{x}}_{ik}$; (ii) subgroup indicators (e.g., same higher-level cluster); (iii) other contextual or structural features relevant to the study. Concrete constructions are illustrated in Section 3 and  4.
@@
initialized from an arbitrary $\mathbf{x}_{(0)}$. This provides a constructive algorithm to recover a valid correlation matrix ${\bf R}=f^{-1}(\boldsymbol{\gamma})$ from an unconstrained vector $\boldsymbol{\gamma}$: (i) embed $\boldsymbol{\gamma}$ into a symmetric matrix ${\bf G}=\operatorname{vecl}^{-1}(\boldsymbol{\gamma})$ with arbitrary diagonal entries; (ii) compute the unique diagonal adjustment $\mathbf{x}^{*}$ using the fixed-point iteration; (iii) obtain the correlation matrix as ${\bf R}=e^{{\bf G}[\mathbf{x}^{*}]}$.
@@
where $C$ is a constant independent of $\boldsymbol{\alpha}$ and $\hat{{\bf V}}_{i}=\hat{{\bf A}}_{i}^{1/2}{\bf R}_{i}(\boldsymbol{\alpha})\hat{{\bf A}}_{i}^{1/2}$. Defining the standardized residual, $\hat{\boldsymbol{\nu}}_{i}=\hat{{\bf A}}_{i}^{-1/2}({\mathbf{y}}_{i}-\hat{\boldsymbol{\mu}}_{i}),$ the pseudo-likelihood estimator of $\boldsymbol{\alpha}_{0}$ can be written as
@@
This approach offers a practical, numerically stable way to estimate correlation parameters in the generalized correlation regression framework. By working in the unconstrained space of the generalized $z$-transformation, it ensures valid correlation matrices while flexibly incorporating covariates. Under standard regularity conditions, the estimator $\hat{\boldsymbol{\alpha}}_{n}$ is consistent for $\boldsymbol{\alpha}_{0}$ and asymptotically normal, providing a sound basis for statistical inference; see the Supplementary Material for technical details. An iterative procedure based on a modified Fisher scoring algorithm can be used to jointly estimate $(\boldsymbol{\alpha},\boldsymbol{\beta},\phi)$, which is also provided in the Supplementary Material.
@@ 3.1 Dataset Overview
3.2 Analysis by Our Correlation Models
We begin by introducing the standardized residuals:
$\hat{\boldsymbol{\varepsilon}}_{i}=(\hat{\varepsilon}_{i1},\dots,\hat{\varepsilon}_{i,m_{i}})^{\top}=\hat{{\bf V}}_{i}^{-1/2}(y_{i1}-\hat{p}_{i1},\dots,y_{i,m_{i}}-\hat{p}_{i,m_{i}})^{\top},$
where $\hat{{\bf V}}_{i}$ is the estimated covariance matrix defined in (3). The empirical correlation for a predefined subgroup $S$ is calculated as:
$\hat{\rho}_{S}=\frac{1}{N_{S}}\sum_{(i,j)\neq(t,s)}\hat{\varepsilon}_{ij}\,\hat{\varepsilon}_{ts}\,\mathbb{I}\big((y_{ij},y_{ts})\in S\big),$ (11)
where $i,t$ index communities, $j,s$ index records, and $N_{S}$ is the number of observation pairs in subgroup $S$. We perform t-tests on each subgroup to examine the significance of the empirical correlations.
Under a correctly specified model, these empirical correlations are expected to be close to zero. This follows from the fact that $\operatorname{Var}(\hat{\boldsymbol{\varepsilon}}_{i})=\hat{{\bf V}}_{i}^{-1/2}{\bf V}_{i}\hat{{\bf V}}_{i}^{-1/2}\approx{\bf I}$ and the assumption of independence between clusters, implying $\mathbb{E}(\hat{\varepsilon}_{ij}\hat{\varepsilon}_{ts})\approx 0$ for $(i,j)\neq(t,s)$. Hence, any statistically significant deviations from zero indicate a lack of fit. Note that when an independent working correlation structure is used, $\hat{{\bf V}}_{i}$ reduces to a diagonal matrix of marginal variances. In this case, $\hat{\rho}_{S}$ can be directly interpreted as an estimate of the correlation of the outcome $y_{ij}$ within subgroup $S$. For other working correlation structures, however, $\hat{\rho}_{S}$ primarily serves as a tool for model diagnostics and variable selection.
A primary source of dependence arises from shared community- and family-level influences, expected to induce correlations among prenatal care choices. To visualize these baseline clustering effects, we define four subgroups and calculate the empirical correlation (11) fitted by the independent GEE. At the community level, let $S_{1}$ denote pairs within the same community ($i=t$ and $(i,j)\neq(t,s)$) and $S_{2}$ its complement. At the family level, $S_{3}$ (within-family) and $S_{4}$ (between-family) are defined analogously. The resulting empirical correlations are shown in the upper panel of Table 2, all results except $\rho_{S_{4}}$ are significant.
Estimate Within-community Between-community Within-family Between-family
E.Corr. $0.1455$ 0.0015-0.0015 $0.7764$ 0.0007-0.0007
pp-Value $0.0000$ $0.0178$ $0.0000$ $0.2934$
E.Corr. $0.0000$ 0.0004-0.0004 $0.0000$ 0.0004-0.0004
pp-Value $0.6527$ $0.5006$ $0.6351$ $0.5206$
Table 2: Empirical correlations for community and family clustering. The upper panel shows results for independent GEE; the lower panel shows results for (12).
Figure 1 presents boxplots of the empirical correlations for the four subgroups, illustrating the presence of strong within-family correlation and moderate within-community correlation. These results indicate that correlations are stronger within communities and within families, with family-level clustering effects notably more pronounced than community-level effects, which are failed to be captured by the independent GEE.
@@ 3.2 Analysis by Our Correlation Models
where $w_{ijk,1}=1$ if births $j$ and $k$ within community $i$ belong to the same family, and 0 otherwise. Under this hierarchical structure, the vector $\boldsymbol{\gamma}_{i}$ exhibits a block pattern:
@@
so that $\alpha_{0}$ captures the baseline within-community correlation, while $\alpha_{1}$ represents the additional within-family correlation. Consequently, the resulting correlation matrix ${\bf R}_{i}$ inherits the same block structure.
@@
The estimated parameters are $\hat{\alpha}_{0}=0.0468$ and $\hat{\alpha}_{1}=0.8617$, both highly significant (p-values <0.0001<0.0001). These results are consistent with the empirical residual correlations and prior findings (Pebley et al.,, 1996; Rodríguez and Goldman,, 2001), confirming strong family-level and moderate community-level clustering effects in the data. Again, we examine the empirical correlations (11) for the four community- and family-level subgroups fitted by our model (12), which are shown in the lower panel of Table 2. It can be seen that all empirical correlations are now insignificant with values approaching zero, suggesting that our proposed model (12) successfully captures the two baseline clustering effects.
Beyond the baseline community- and family-level clustering, other features may induce additional correlations. For example, the ethnic-linguistic identifier indig links multiple birth records, serving as a natural ethnic-linguistic grouping variable.
We begin the exploration by computing the empirical correlations (11) within subgroups defined by family- and individual-level covariates. For ethnic-linguistic identities, three subgroups are defined: a pair of births is considered to belong to the same subgroup only if they share both the same community and the same ethnic-linguistic identity, which corresponds to the three ethnic groups. A similar approach is applied to husband’s employment status. For other family- and individual-level covariates, subgroups are constructed based on absolute differences. For example, in the case of mother’s education level, three subgroups are defined according to whether the absolute difference in education between a birth pair is 0, 1, or 2, while again restricting to pairs within the same community. The empirical correlations are fitted via our initial correlation model (12). While other covariates tend out to be insignificant, two covariates stand out—ethnic-linguistic identity and husband’s employment status, with results shown in the upper panel of Table 3 and 4.
Estimate Ladino Indigenous non-Spanish Indigenous Spanish
E.Corr. $0.0040$ $0.0230$ 0.0362-0.0362
pp-Value $0.7146$ $0.3428$ $0.0817$
E.Corr. $0.0029$ $0.0131$ 0.0047-0.0047
pp-Value $0.7905$ $0.5996$ $0.8120$
Table 3: Empirical correlations for ethnic-linguistic identity. The upper panel shows results for (12); the lower panel shows results for (13).
Estimate Unskilled Professional Agri-self Agri-empl Skilled
E.Corr. $0.1002$ 0.0168-0.0168 $0.0509$ $0.0057$ $0.0066$
pp-Value $0.2947$ $0.7225$ $0.0131$ $0.8292$ $0.8045$
E.Corr. 0.1372-0.1372 0.0409-0.0409 $0.0209$ $0.0115$ 0.0266-0.0266
pp-Value $0.4936$ $0.4115$ $0.3092$ $0.6484$ $0.3255$
Table 4: Empirical correlations for husband’s employment status. The upper panel shows results for (12); the lower panel shows results for (13).
The empirical correlations are marginally significant (p <0.1<0.1) for the indigenous Spanish subgroup and significant (p <0.05<0.05) for the agri-self subgroup. This indicates that our initial correlation model (12) is insufficient to capture the additional correlation patterns beyond the simple two-level hierarchical clustering within these subgroups, thus demanding the need to extend our correlation model by incorporating the two covariates indig (ethnic-linguistic identity) and husEmpl (husband’s employment status).
To illustrate the flexibility and practical utility of our proposed approach, we incorporate the two covariates accordingly, leading to the following correlation model:
$\gamma_{ijk}=$ $\ \alpha_{0}+\alpha_{1}\,w_{ijk,1}+\sum_{l=0}^{2}\alpha_{2+l}\,w_{ijk,2+l}+\sum_{l=0}^{4}\alpha_{5+l}\,w_{ijk,5+l},$ (13)
where the indices $i,j,k$ follow the same convention as in model (12). The additional correlation covariates are defined as:
  • $w_{ijk,5+l}=\mathbb{I}(\text{husEmpl}_{ij}=\text{husEmpl}_{ik}=l)$, for $l=0,1,2,3,4$, indicating shared employment status of the husband.
Beyond generic community- and family-level clustering, model (13) allows correlation to vary with shared ethnic-linguistic background, socioeconomic similarity—factors relevant to prenatal care access and choices. The parameter estimates and corresponding $p$-values for model (13) are summarized in Table 5.
Variable Estimate P-value
Intercept ($\alpha_{0}$) 0.05410.0541* $0.0459$
mom ($\alpha_{1}$) 0.81900.8190*** $0.0000$
Ladino ($\alpha_{2}$) 0.0195-0.0195 $0.5512$
Indigenous non-Spanish ($\alpha_{3}$) 0.0141-0.0141 $0.7405$
Indigenous Spanish ($\alpha_{4}$) 0.0626-0.0626. $0.0684$
Unskilled ($\alpha_{5}$) $0.3635$ $0.2859$
Professional ($\alpha_{6}$) $0.0761$ $0.7812$
Agri-self ($\alpha_{7}$) $0.0323$. $0.0881$
Agri-empl ($\alpha_{8}$) $0.0122$ $0.6439$
Skilled ($\alpha_{9}$) $0.0653$ $0.2667$
Notes: . for $\text{p}<0.1$, * for $\text{p}<0.05$, ** for $\text{p}<0.01$ and *** for $\text{p}<0.001$.
Table 5: Parameter estimates and corresponding p-values for (13).
The estimation results aligns well with the empirical studies above. In model (13), the hierarchical clustering effects $\hat{\alpha}_{0}$ and $\hat{\alpha}_{1}$ remain significant, while most parameters are insignificant, likely because the strong baseline clustering captures most of their potential variation. Two exceptions are again $\hat{\alpha}_{4}$ (indigenous Spanish ethnic-linguistic group) and $\hat{\alpha}_{7}$ (husband agri-self employment), both marginally significant (p <0.1<0.1), indicating additional dependence beyond generic family- and community-level effects. Moreover, the empirical correlations (11) for the baseline clustering and all the covariate-based subgroups fitted via our model (13) are now no longer significant. Specifically, the empirical results for ethnic-linguistic identity and husband’s employment status are shown in the lower panel of Table 3 and 4. This suggests that our model (13) has already adequately capture all the dependencies associated with family and individual level covariates.
Even after accounting for general family and community clustering, births in the indigenous Spanish population are somewhat less correlated (negative estimated value), whereas mothers in agri-self households show greater similarity in prenatal care choices (positive estimated value). The finding of weaker within-group correlation among the indigenous Spanish population may be attributed to their higher degree of cultural assimilation, which could promote individual behavioral diversity and reduce uniform group practices. The stronger correlation observed within agriculturally self-employed households might be explained by occupational isolation. This isolation can foster socioeconomic homogeneity and shared experiences, thereby increasing behavioral consistency in healthcare decisions. These patterns reflect cultural norms or occupation-related constraints, suggesting potential targets for health interventions or further investigation.
3.3 Comparisons with GLMM
For computational efficiency, we consider a simplified version of model (13):
$\gamma_{ijk}=$ $\ \alpha_{0}+\alpha_{1}\,w_{ijk,1}+\alpha_{2}\,w_{ijk,2}+\alpha_{3}\,w_{ijk,3}+\alpha_{4}\,w_{ijk,4}+\alpha_{5}\,w_{ijk,5},$ (14)
where the indices $i,j,k$ and covariate $w_{ijk,1}$ are the same as in model (12). The additional correlation covariates are defined as:
  • $w_{ijk,2}=\mathbb{I}(\text{indig}_{ij}=\text{indig}_{ik}\neq 2)$ and $w_{ijk,3}=\mathbb{I}(\text{indig}_{ij}=\text{indig}_{ik}=2)$, indicating shared ethnic-linguistic identity with the indigenous Spanish taken out separately;
  • $w_{ijk,4}=\mathbb{I}(\text{husEmpl}_{ij}=\text{husEmpl}_{ik}\neq 2)$ and $w_{ijk,5}=\mathbb{I}(\text{husEmpl}_{ij}=\text{husEmpl}_{ik}=2)$, indicating shared employment status of the husband with the agri-self taken out separately.
Fitting the simplified model (14) obtains similar results as in model (13): $\alpha_{0}$ and $\alpha_{1}$ are significantly positive with $\alpha_{1}$ having the larger absolute value; $\alpha_{3}$ are $\alpha_{5}$ are marginally significant with the same signs as in model (13), while $\alpha_{2}$ are $\alpha_{4}$ are insignificant.